The expression of the electric field associated with an electromagnetic wave in vacuum is given by



E vector =(200Vm^-1) x^ . sin(2π×10^8 t -ky)


Determine the the direction of propagation, wave number, frequency, and the


magnitude and direction of the magnetic field associated with the wave.



1
Expert's answer
2022-02-28T13:37:14-0500

Answer

In +x direction.

Wave no. is given

=k

Frequecy

"=2*3.14*10^8\\=4.28*10^8rad/sec"

Magnitude

=200V/m

In y direction.